OpenStudy (anonymous):

Help!!! k-3/4=2/5. How do I do this?

OpenStudy (anonymous):

cross multiply

OpenStudy (anonymous):

so its 5K -15= 8

OpenStudy (anonymous):

\[k=(3/4)+(2/5)\] \[k=(15/20)+(8/20)\] \[k=23/20\]

OpenStudy (anonymous):

wait is it k-3 divided by 4?

OpenStudy (anonymous):

k= 23/5

OpenStudy (anonymous):

k - (3/4) = (2/5)

OpenStudy (anonymous):

oh ok should of made that more clear, sorry. then yea. ^ the other guy is right.

OpenStudy (anonymous):

sorry - thanks all for the help. may have more.

OpenStudy (anonymous):

ok cool. let me know i'll see if i can help.

OpenStudy (anonymous):

f + l-3(11/12)l=18